### v1

@@ 2.3 AI Bias and Disparities for ELLs
Bias and uneven treatment of ELLs and non-ELLs have always been a concern in traditional educational environments. For instance, science teachers in the US, who are mostly English speakers, can view ELLs through a deficit lens and thus expect less from them and assess them according to those lower expectations [17]. With the introduction of AI in science education, it is not surprising that educators and researchers have expressed concerns about whether the use of AI may exacerbate the existing bias in education that mostly impacts minoritized ELL groups [2]. Theoretically, the AI scoring model could have a statistical and computational bias if the training dataset lacks representation of certain groups, since the AI scoring model would reflect the pattern of the training dataset [21]. When it comes to ELLs, although the proportion of ELL students is increasing rapidly in some multicultural and multilingual countries like the US, the percentage of ELLs compared to non-ELLs is significantly much lower (e.g., 1:9 in the US), which would result in unbalanced datasets for AI model training. Meanwhile, the characteristics of linguistic features of ELLs and non-ELLs are obvious, which makes it a potential pattern to influence AI model scoring on ELLs’ responses. AI may struggle to interpret nonstandard or unexpected responses of ELLs, resulting in misinterpretation or underestimation of valid yet diverse expressions of knowledge from students with different English proficiency levels  cite cheuk2021can. This situation can perpetuate existing inequities in educational assessment.
@@ 5 Discussion and Conclusion
Our findings are unsurprising and contribute to the ease of unnecessary fears of AI bias, which are critical for all populations to benefit from AI[27]. Altough many studies have found automatic scoring bias for ELLs, similar to this study, Wilson and Huang (2024) found that when using an automatic scoring system for writing tests, the system did not exhibit unique bias against ELLs but rather replicated biases present in human scoring [24]. Reed and Mercer (2022) also found no evidence of bias for ELLs when comparing interim assessments scored by teachers or expert raters to predict student performance on a summative writing test [20]. Meanwhile, empirical studies that found English proficiency to be an important factor that impacts the automatic scoring or categorization results of student writing tasks, are primarily focused on essay writing tasks [18] [20] [24]. These tasks asked students to write quite long responses, and their writing proficiency is often a key evaluation criterion, which is highly related to their English proficiency. Unlike essay writing assessments, science constructed-response items are typically shorter and less dependent on language proficiency, reducing the impact of language proficiency on AI scoring. Moreover, science assessments may include multimodal components, such as diagrams, tables, and symbols, which support students’ understanding of the concepts in the question and help the accurate interpretations of the students’ understanding beyond their language proficiency [6] [7].Thus, the automatic scoring of these responses is less likely to be influenced by students’ English proficiency compared to previous research on essay writing tasks. Our findings therefore suggest that AI scoring neither amplifies existing bias nor distorts true disparities for ELLs.
@@ 6 Limitation and Future Study
Although the results of this study show that AI scoring generally does not lead to biased results for ELLs and non-ELLs in science constructed response assessments, it is important to note that AI bias does not solely stem from the factor investigated (i.e., the training dataset in this study). Schwartz et al. (2022) [21] classify AI biases into three groups: systemic, statistical, computational, and human biases. Systemic biases, such as sexism and racism, are inherited in datasets. They arise from the historical and institutional norms, as well as societal practices that privilege certain groups of people. Statistical and computational biases are caused by the lack of representation or misrepresentation of certain groups during AI training. Human biases, on the other hand, are mostly related to the user and AI interaction, such as biases in prompts, interpretation, confirmation, and the use of the outputs. Therefore, future studies should continue to address the potential AI bias for ELL students, considering all these possible dimensions. Moreover, future research can provide an even more nuanced approach by exploring whether AI scoring bias and disparities exist for different levels of English proficiency, such as WIDA’s proficiency level distinctions, rather than a broad ELL/non-ELL binary [23].

### v2

@@ Abstract
Cite the study: Guo, S., Wang, Y., Yu, J. Wu, X., Ayik, B., Watts, F., Latif, E., Liu, L., Liu, N., & Zhai, X. (2025). Artificial Intelligence Bias on English Language Learners in Automatic Scoring. Proceedings of the International Conference on AI in Education, pp. 1-12, Palermo, Italy.
@@ 2.3 AI Bias and Disparities for ELLs
Bias and uneven treatment of ELLs and non-ELLs have always been a concern in traditional educational environments. For instance, science teachers in the US, who are mostly English speakers, can view ELLs through a deficit lens and thus expect less from them and assess them according to those lower expectations [17]. With the introduction of AI in science education, it is not surprising that educators and researchers have expressed concerns about whether the use of AI may exacerbate the existing bias in education that mostly impacts minoritized ELL groups [2]. Theoretically, the AI scoring model could have a statistical and computational bias if the training dataset lacks representation of certain groups, since the AI scoring model would reflect the pattern of the training dataset [21]. When it comes to ELLs, although the proportion of ELL students is increasing rapidly in some multicultural and multilingual countries like the US, the percentage of ELLs compared to non-ELLs is significantly much lower (e.g., 1:9 in the US), which would result in unbalanced datasets for AI model training. Meanwhile, the characteristics of linguistic features of ELLs and non-ELLs are obvious, which makes it a potential pattern to influence AI model scoring on ELLs’ responses. AI may struggle to interpret nonstandard or unexpected responses of ELLs, resulting in misinterpretation or underestimation of valid yet diverse expressions of knowledge from students with different English proficiency levels [2]. This situation can perpetuate existing inequities in educational assessment.
@@ 5 Discussion and Conclusion
Our findings are unsurprising and contribute to the ease of unnecessary fears of AI bias, which are critical for all populations to benefit from AI[27]. Although many studies have found automatic scoring bias for ELLs, similar to this study, Wilson and Huang (2024) found that when using an automatic scoring system for writing tests, the system did not exhibit unique bias against ELLs but rather replicated biases present in human scoring [24]. Reed and Mercer (2022) also found no evidence of bias for ELLs when comparing interim assessments scored by teachers or expert raters to predict student performance on a summative writing test [20]. Meanwhile, empirical studies that found English proficiency to be an important factor that impacts the automatic scoring or categorization results of student writing tasks, are primarily focused on essay writing tasks [18] [20] [24]. These tasks asked students to write quite long responses, and their writing proficiency is often a key evaluation criterion, which is highly related to their English proficiency. Unlike essay writing assessments, science constructed-response items are typically shorter and less dependent on language proficiency, reducing the impact of language proficiency on AI scoring. Moreover, science assessments may include multimodal components, such as diagrams, tables, and symbols, which support students’ understanding of the concepts in the question and help the accurate interpretations of the students’ understanding beyond their language proficiency [6] [7].Thus, the automatic scoring of these responses is less likely to be influenced by students’ English proficiency compared to previous research on essay writing tasks. Our findings therefore suggest that AI scoring neither amplifies existing bias nor distorts true disparities for ELLs.
@@ 6 Limitation and Future Study
Although the results of this study show that AI scoring generally does not lead to biased results for ELLs and non-ELLs in science constructed response assessments, it is important to note that AI bias does not solely stem from the factor investigated (i.e., the training dataset in this study). Schwartz et al. [21] classify AI biases into three groups: systemic, statistical, computational, and human biases. Systemic biases, such as sexism and racism, are inherited in datasets. They arise from the historical and institutional norms, as well as societal practices that privilege certain groups of people. Statistical and computational biases are caused by the lack of representation or misrepresentation of certain groups during AI training. Human biases, on the other hand, are mostly related to the user and AI interaction, such as biases in prompts, interpretation, confirmation, and the use of the outputs. Therefore, future studies should continue to address the potential AI bias for ELL students, considering all these possible dimensions. Moreover, future research can provide an even more nuanced approach by exploring whether AI scoring bias and disparities exist for different levels of English proficiency, such as WIDA’s proficiency level distinctions, rather than a broad ELL/non-ELL binary [23].